A small ball is suspended from a thread. It is lifted up with an acceleration 4.9 ms(sq) and lowered with an acceleration 4.9ms(sq) then the ratio of tensions in the thread in both cases will be :-
1
Expert's answer
2019-04-12T09:40:31-0400

Earth gravitational acceleration: "g=9.8 \, \text{ms}^{-2}"

When a small ball is lifted up with an acceleration "a=4.9 \, \text{ms}^{-2}" , the tension of the thread equals to


"T_u=m(g+a)"

When it is lowered with the same acceleration, , the tension of the thread equals to

"T_d=m(g-a)"

The ratio:


"\frac{T_u}{T_d} = \frac{g+a}{g-a}=\frac{9.8+4.9}{9.8-4.9}=3"